OpenStudy (anonymous):

Find the intersection point(s) of the graphs of y=x^2 and 6+y=5x . Plot the intersection points (not the graphs of the equations) on the axes below.

OpenStudy (mertsj):

y=5x-6 from second equation. Equate the y's x^2=5x-6 x^2-5x+6=0 (x-3)(x-2)=0 x=3 and x=2 are the x coordinates of the points of intersection. So plug 3 and 2 into either equation to get y. The graphs intersect at (3,9) and (2,4)

OpenStudy (mertsj):